OpenStudy (anonymous):

What values for theta(0< or equal to theta < or equal to 2pi) satisfy the equation? 3sin theta = sin theta - 1 tan^2 theta = -3/2sec theta Help! I have no idea what im doing

OpenStudy (campbell_st):

for the 1st equation \[3\sin(\theta) = \sin(\theta) - 1\] so start by subtracting \[\sin(\theta)\] from both sides of the equation... what do you get..?

OpenStudy (anonymous):

ummm 4 sin=-1?

OpenStudy (anonymous):

or 2sin=-1?

OpenStudy (campbell_st):

its the 2nd option \[2 \sin(\theta) =-1\] if you divide both sides of the equation by 2 you get \[\sin(\theta) = -\frac{1}{2}\] does that make sense..?

OpenStudy (anonymous):

yes... so 2 would be tan^2 = -3/2 sec sin/cos ^2 = -3/2 1/cos = sin^2/cos=-3/2 ?

OpenStudy (campbell_st):

well getting back to the 1st equation... you can now solve this... find sin = 1/2 and sin is negative in the 3rd and 4th quadrants so the answer is 180 + theta and 360 - theta hope that is ok..

OpenStudy (anonymous):

pi/3 and 5pi/3

OpenStudy (anonymous):

??

OpenStudy (campbell_st):

for the 2nd equation rewrite it as \[\frac{\sin^2(\theta)}{\cos^2(\theta)} = -\frac{3}{2\cos(\theta)}\] multiply both sides by \[\cos^2(\theta)\] and you get \[\sin^2(\theta) = \frac{-3\cos(\theta)}{2}\] then using sin^2 + cos^2 = 1 you can make the substitution \[1 - \cos^2(\theta) = -\frac{3\cos(\theta)}{2} \] which becomes \[\cos^2(\theta) - \frac{3\cos(\theta)}{2} - 1 = 0\] so you can solve the quadratic hope it makes sense

OpenStudy (campbell_st):

ok... so if you are working in radians the angles are \[\pi + \theta~~and~~~2\pi - \theta\]

OpenStudy (anonymous):

Now you lost me completely.

OpenStudy (campbell_st):

remember from an exact value triangle \[\sin(\theta) = \frac{1}{2}~~~then~~~~\theta = \frac{\pi}{6}\]

OpenStudy (anonymous):

But there must also be a second one... so pi/6 and 3pi/6?

OpenStudy (campbell_st):

here is the diagram  since sin is negative it means the angles are in the 3rd and 4th quadrants... so the 3rd quadrant angle is \[\pi + \frac{\pi}{6}\] and the 4th quadrant angle is \[2\pi - \theta\]